OpenStudy (anonymous):

Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. A = 61°, a = 17, b = 19

OpenStudy (anonymous):

Two triangles can not be formed with the given information. The information gives only the lengths of 2 legs and their angle of intersect.

OpenStudy (anonymous):

BTW, the 2 legs can be closed, and the triangle "solved" using sines, but improper questions make me antsy.

OpenStudy (anonymous):

\[\frac{\sin(B)}{19}=\frac{\sin(61)}{17}\] \[B=\arcsin(\frac{19\sin(61)}{17})\]

OpenStudy (anonymous):

now you have to angles, 61 and 77.83 you can easily find the third, since the sum is 180

OpenStudy (anonymous):

*two

OpenStudy (anonymous):

if there is another one, the angle B will be \(180-77.83=102.17\) .

OpenStudy (anonymous):